# Algebraic integers with continued fraction expansions containing palindromes and square roots with prescribed periods

### v1

@@ Abstract.
We prove that there exist infinitely many algebraic integers with continued fraction expansion of the kind $[a_{0},\overline{a_{1},\ldots,a_{n},s}]$ where $(a_{1},\ldots,a_{n})$ is a palindrome and $s\in\mathbb{N}_{\geq 1}$, characterizing all the algebraic integers with such expansions. We also provide an explicit method for finding $s$ and determining the corresponding algebraic integer. Moreover, we deal with the particular case $(a_{1},\ldots,a_{n})=(m,\ldots,m)$ describing the corresponding algebraic integers in terms of Fibonacci polynomials. We exploit these results for obtaining expansions of square roots of integers with prescribed periods and we also write explicitly the fundamental solutions of the corresponding Pell’s equations.
@@ 1. Introduction
In this paper, we firstly address similar problems for algebraic integers $\alpha$ such that $10<\alpha<$ 1. In particular we prove the following results.
Theorem A.
  • (a)
    Given a finite palindromic sequence $(a_{1},\ldots,a_{n})$ there exist infinitely many algebraic integers $\alpha$ of degree two such that $10<\alpha<$ 1 and $\alpha=[0,\overline{a_{1},\ldots,a_{n},s}]$
  • (b)
    Given $\alpha=[0,\overline{m,\ldots,m,s}]$ there exist infinitely many values of $s$ such that $\alpha$ is an algebraic integer
Part (a) deals with the analogue problem solved by Friesen [2] and Mc Laughlin [7]. The proof is given in Section 2 (Theorem 1) where we also provide explicitly the minimal polynomial of such algebraic integers and we give explicit conditions for obtaining all the possible values of $s$. Furthermore, Corollary 1 points out that Theorem 1 characterizes also the algebraic integers $\alpha\not\in(0,1)$ with such expansions. Part (b) considers the analogue situation studied in [12] for square roots. The proof is given in Section 3 (Theorem 8) where the minimal polynomial of such algebraic integers is described in terms of Fibonacci polynomials. Finally, we find infinitely many families of square roots of integers with prescribed periods (Theorems 9 and 10, Remark 11) as well as providing explicit forms for the fundamental solutions of the corresponding Pell’s equations (Section 4).
@@ 2. Algebraic integers with palindromic sequences in their continued fraction expansion
In the following theorem, we prove that for any palindromic sequence there exist infinitely many algebraic integers containing such sequence in the period of the continued fraction expansion. Moreover, in the proof, we explicitly provide the minimal polynomial of these algebraic integers so that they are effectively given.
@@
The remaining cases for the expansion of $\beta(n,m,k)$ can be proved similarly to Theorems 9 and 10 and the resulting expansions are slightly different. Observe that it is convenient to consider the different cases $m=1$ and $m\geq 3$ odd combined with the possible values of $n$ modulo 6. This is due to the fact that the parity of $s$ and $k$ changes depending on these values for $m$ and $n$.

### v2

@@ Abstract.
We present a characterization of the algebraic integers with continued fraction expansions of the form $[a_{0},\overline{a_{1},\ldots,a_{n},s}]$, where $(a_{1},\ldots,a_{n})$ is a palindrome and $s\in\mathbb{N}_{\geq 1}$. In particular, we focus on the special case where $(a_{1},\ldots,a_{n})=(m,\ldots,m)$, providing a detailed characterizations of the corresponding algebraic integers and $s$ in terms of Fibonacci polynomials. Then, we derive new expansions of square roots of integers with these periods, given $m$ and $n$. Moreover, we explicitly determine the fundamental solutions of both positive and negative Pell’s equations corresponding to this family of integers.
2020 Mathematics Subject Classification:
11A55
@@ 1. Introduction
The paper is structured as follows. In Section 2, we give a characterization of all quadratic algebraic integers having continued fraction expansion $[a_{0},\overline{a_{1},\ldots,a_{n},s}]$ where $(a_{1},\ldots,a_{n})$ is a palindrome and $s\in\mathbb{N}_{\geq 1}$, in Theorem 2 and Corollary 1. We also provide an explicit method for finding $s$ and determining the corresponding algebraic integer. This result is not new (since it can be derived from [7, Theorem 3.1], [2] and [8]), however, we include it alongside our proof (slightly different from the approach that exploits [7, Theorem 3.1]) as it constitutes a fundamental component in the subsequent findings.
In Section 3, in Theorem 9, we specify this study to the special case $(a_{1},\ldots,a_{n})=(m,\ldots,m)$, describing the minimal polynomial of such algebraic integers in terms of Fibonacci polynomials. Moreover, we find the explicit continued fraction expansion of a infinite family of square roots of integers (Theorems 10 and 11, Remark 12), derived from algebraic integers of Theorem 9.
In Section 4 we provide the explicit forms for the fundamental solutions of both positive and negative Pell’s equations corresponding to these integers.
@@ 2. Algebraic integers with palindromic sequences in their continued fraction expansion
In the following theorem, we prove that for any palindromic sequence there exist infinitely many algebraic integers containing such sequence in the period of the continued fraction expansion.
Remark 1.
We point out that Theorem 2 is readily deducible from [7, Theorem 3.1]. However, we provide our own proof as it explicitly provides the continued fraction expansion and the minimal polynomial of the algebraic integers in the interval $(0,1)$ so that they are effectively given. These results are particularly pertinent to Theorem 9, where we examine the specific case of a palindrome sequence composed of constant elements.
@@
The remaining cases for the expansion of $\beta(n,m,k)$ can be proved similarly to Theorems 10 and 11 and the resulting expansions are slightly different. Observe that it is convenient to consider the different cases $m=1$ and $m\geq 3$ odd combined with the possible values of $n$ modulo 6. This is due to the fact that the parity of $s$ and $k$ changes depending on these values for $m$ and $n$.